Has Gold Lost Its Luster?

Please use ONLY one Excel file to complete the case study and upload the Excel file to the submission link (Week 3 Case Study) for grading. Please clearly mark your answers (either highlight or font colors) and grammar counts.

No credit will be granted for problems that are not completed using Excel.

In 2011, when the Gallup organization polled investors, 34% rated gold the best long-term investment. In April of 2013 Gallup surveyed a random sample of U.S. adults. Respondents were asked to select the best long-term investment from a list of possibilities. Only 241 of the 1005 respondents chose gold as the best long-term investment. By contrast, only 91 chose bonds.

  1. Compute the standard error for each sample proportion. Compute and describe a 95% confidence interval in the context of the question.
  2. Do you think opinions about the value of gold as a long-term investment have really changed from the old 34% favorability rate, or do you think this is just sample variability? Explain.
Answer #1

a) standard error of proportion = sqrt(pq/n)

for gold

get sd(p^)
n 1005
p 0.3400
sd(p^) 0.014943

Formula

get sd (p^)
n 1005
p 0.34
sd(p^) =SQRT(B3*(1-B3)/B2)

95% confidence interval

For Gold

get CI
X 241
n 1005
p 0.2398
q 0.7602
alpha 0.05
z 1.9600
E 0.0264
lower 0.2134
upper 0.2662

(0.2134, 0.2662)

Formula

get CI
X 241
n 1005
p =B2/B3
q =1-B5
alpha 0.05
z =NORMSINV(1-B8/2)
E =B9*SQRT(B6*B5/B3)
lower =B5-B11
upper =B5+B11

For bonds

get CI
X 91
n 1005
p 0.0905
q 0.9095
alpha 0.05
z 1.9600
E 0.0177
lower 0.0728
upper 0.1083

(0.0728, 0.1083)

b) since 0.34 is not present in confidence interval (0.2134, 0.2662)

hence opinions about the value of gold as a long-term investment have really changed from the old 34% favorability rate